A car, initially at rest, travels from 0 m/s to 15.2 m/s in 20.3 s. What is the car's acceleration?


1
Expert's answer
2021-04-12T07:04:12-0400

The acceleration is

"\displaystyle a = \frac{dv}{dt} = \frac{\Delta v}{\Delta t} = \frac{v_2 - v_1}{t} = \frac{15.2-0}{20.3}= 0.7487 \approx 0.75 \; m/s^2"


Answer: "0.75\;m/s^2"